OpenStudy (anonymous):

Find the exact value by using a half-angle identity. tan7π/8 PLEASE PLEASE PLEASE :)

OpenStudy (anonymous):

Well, the half angle identity says: \[ \sin(\theta/2) = \sqrt{\frac{1-\cos(\theta)}{2}} \quad \cos(\theta/2) = \sqrt{\frac{1+\cos(\theta)}{2}} \]

OpenStudy (anonymous):

but what about tan..

OpenStudy (anonymous):

Well \[ \tan(\theta)=\frac{\sin(\theta)}{\cos(\theta)} \]So I didn't really remember it's identity for this. Rather, I just divided the other identities.

OpenStudy (anonymous):

So then \[ \tan(\theta / 2) = \sqrt{\frac{1-\cos(\theta)}{1+\cos(\theta)}} \]

OpenStudy (anonymous):

well i just don't really understand how to do the problem once everything is plugged in

OpenStudy (anonymous):

Well, \(\cos(7\pi/4)\) is relatively easy to solve for, when you look at the unit circle.

OpenStudy (anonymous):

well i know how to find the degree but once i get the degree i dont know where to go from there

OpenStudy (anonymous):

that is a good method... but tan's identity is below \[\large \tan (\frac{ \theta }{ 2 }) =\frac{ \sin(\theta) }{1+ \cos (\theta) }=\frac{ 1- \cos (\theta) }{\sin(\theta) }\]

OpenStudy (anonymous):

do you mind just showing me the steps please??

OpenStudy (anonymous):